### v2

# Łukasz Struski 1,  Dawid Rymarczyk 1,2,  Jacek Tabor 1

Abstract
In this work, we introduce InfoDisent, a hybrid approach to explainability based on the information bottleneck principle. InfoDisent enables the disentanglement of information in the final layer of any pretrained model into atomic concepts, which can be interpreted as prototypical parts. This approach merges the flexibility of post-hoc methods with the concept-level modeling capabilities of self-explainable neural networks, such as ProtoPNets. We demonstrate the effectiveness of InfoDisent through computational experiments and user studies across various datasets using modern backbones such as ViTs and convolutional networks. Notably, InfoDisent generalizes the prototypical parts approach to novel domains (ImageNet).
Introduction
Deep neural networks have demonstrated performance that matches or even surpasses human capabilities across various domains, such as image classification and generation, speech recognition, and natural language processing. Despite their impressive achievements, these networks often operate as ”black boxes”, offering little insight into the reasoning behind their decisions (rudin2019stop).
This lack of transparency poses significant challenges, particularly in high-stakes applications such as medicine and autonomous driving, where understanding a model’s decision-making process is critical (bojarski2017explaining; khan2001classification; nauta2023anecdotal; patricio2023explainable; samek2021explaining; struski2024multiple). To address this issue, the subfield of artificial intelligence known as eXplainable AI (XAI) (xu2019explainable) has emerged, focusing on making AI systems more interpretable and trustworthy.
XAI methods can be classified into two major categories: post-hoc methods and inherently interpretable (ante-hoc) methods (rudin2019stop). Post-hoc methods, such as GradCAM (selvaraju2017grad), are highly flexible because they can be applied to any neural network architecture. However, they are often unreliable (adebayo2018sanity), provide local explanations, and fail to reveal the characteristics of specific classes. In contrast, ante-hoc methods, such as ProtoPNet (chen2019looks), offer concept-level explanations by identifying prototypical parts from the training dataset. While this approach enhances interpretability, it is
Figure 1: Decision explanation constructed by InfoDisent for the pre-trained ViT feature space on the Agaric mushrooms image from the ImageNet. We can trace the decision of ViT behind assigning the class Agaric to the image on the left to having a hat (569), a white leg (728), a reddish shine (552), a strawberry texture (297) and the appearance of ground with moss (311). Note that, the prototype block (right) each row represents the prototypical part (the corresponding channel number). The yellow boxes in each row show the activation of a given prototypical part, while in the first column, we show the activation of corresponding prototypical parts in the original image.
limited to fine-grained classification tasks and specific architectures, such as CNNs in the case of ProtoPNet. Consequently, these methods do not fully leverage the potential of large pretrained models, such as Vision Transformers (ViTs) (dosovitskiy2020image).
To address this, we propose InfoDisent, a novel XAI approach that leverages information disentanglement in the final layer of any pretrained model. In InfoDisent, each channel encapsulates a single atomic concept, interpretable as prototypical parts – similar to the approach used in PIPNet (nauta2023pipnet). This is achieved through the application of an information bottleneck, in which we enforce activation sparsity in each prototypical channel. As a result, InfoDisent provides both local and global-level explanations in the form of human-interpretable concepts, as illustrated in Figure 1. Furthermore, InfoDisent offers significant flexibility, as it can be seamlessly applied to any pretrained architecture, including ViTs and convolutional networks.
We demonstrate the effectiveness of InfoDisent through both computational experiments and user studies. Our method is benchmarked on five datasets, including ImageNet, a challenging dataset where prototypical parts-based approaches have not been generalized to. Additionally, results from the user study indicate that InfoDisent provides a competitive level of explanation understanding compared to other prototypical parts-based methods, while offering the advantage of flexibility in its application to any model backbone. We make the code available.
Our contributions can be summarized as follows:
  • We propose InfoDisent, a hybrid XAI model that combines the interpretability of ante-hoc methods with the flexibility of post-hoc approaches.
  • To the best of our knowledge, InfoDisent is the first model to provide prototypical, part-like interpretations for the feature space of any pretrained network.
  • We validate the effectiveness of InfoDisent in terms of both accuracy and user understanding through extensive experimental evaluations.
Related Works
Research in XAI can be divided into two disjoint categories: post-hoc interpretability (lundberg2017unified; ribeiro2016should; selvaraju2017grad), where we analyze the pre-trained model to explain its predictions, and inherently explained models (bohle2022b; chen2019looks), where the aim lies in building networks which decisions are easy to interpret. Both of the above approaches have their advantages and disadvantages which we discuss in the following paragraphs.
Post-hoc methods
In the post-hoc methods, we interpret existing pre-trained network architectures. The commonly used methods such as SHAP (lundberg2017unified; shapley1951notes), LIME (ribeiro2016should), LRP (bach2015pixel) and Grad-CAM (selvaraju2017grad) provide in practice only feature importance which can be visualized as a saliency map that shows on which part of the image the model has focused its attention. This allows us to check if the model does not focus its attention outside of the object of interest (ribeiro2016should), however, it is in general not sufficient to really understand the reasons behind given predictions. Additionally, post-hoc methods allow typically only local explanations (per the prediction of a given image), and do not allow to understand of the prerequisites to the given class.
Inherently explained models
While post-hoc methods are easy to implement due to their non-intrusive nature, they often produce biased and unreliable explanations (NEURIPS2018_294a8ed2). To address this, recent research has increasingly focused on designing self-explainable models that make the decision process directly visible (brendel2018approximating; NEURIPS2018_3e9f0fc9). Many of these interpretable solutions utilize attention mechanisms (liu2021visual; zheng2019looking) or exploit the activation space, such as with adversarial autoencoders (Guidotti_Monreale_Matwin_Pedreschi_2020). Among the most recent approaches, ProtoPNet (chen2019looks) has significantly influenced the development of self-explainable models. It learns class-specific prototypes, similar to concepts, with a fixed number per class. The model classifies inputs by calculating responses from each class’s prototypes and summarizing these responses through a fully connected layer, providing explanations as a weighted sum of all prototypes. This method inspired the development of several other self-explainable models (donnelly2022deformable; nauta2023pipnet; pach2024lucidppn; rymarczyk2021protopshare; rymarczyk2022interpretable; rymarczyk2023icicle; wang2021interpretable). Typically, in the prototypical parts-based model, the final decision of a given class is decomposed into the appearance of a few selected prototypes, which are similar to some strongly localized parts of some chosen images from the training dataset. We see that in this approach the final class decision is split into a set of simpler and interpretable concepts.
Figure 2: Our image classification interpretation model, InfoDisent, features three main components: a pre-trained backbone, a pooling layer for key features, and a fully connected layer. The CNN/transformer backbone, with frozen weights, is not further trained. The pooling layer extracts features from the last transformer or convolutional layer and identifies key positive and negative features. These are then combined into a dense vector, which is processed by a fully connected linear layer with nonnegative entries in the final stage.
InfoDisent
Our motivation comes from the general methodology of prototypical models, which can be described as follows:
  • trace the final class decision to the co-occurrence in the image of some prototypes,
  • each prototype can be visualized by some prototypes from the training dataset,
  • prototypes should be strongly localized parts of the image.
We claim that even for pre-trained models it is possible to disentangle the channels in the feature space in such a way that they will become sufficiently informative to satisfy the above points. In InfoDisent we disentangle the channels in feature space by applying an orthogonal map in the pixel space, and consequently do not change the inner-lying distance and scalar product.
Since we disentangle the channels in the feature space, we restrict our attention to images in the feature space and consider only the classification head. Thus consider the data set of images in the feature space, which consists of images of possibly different resolutions but the same number of channels $d$.
Default classification head in deep networks
To establish notation, let us first describe the typical classification head. In the case of a classification task with $k$ classes, we apply the following operations for the image $I$ in the feature space:
  1. 1.
    $I\to v_{I}=\mathrm{avg\_pool\_over\_channels}(I)\in\mathbb{R}^{d}$
  2. 2.
    $v_{I}\to w_{I}=Av_{I}$, where $A$ is a matrix of dimensions $d\times k$
  3. 3.
    $w_{I}\to p_{I}=\mathrm{softmax}(w_{I})$.
In the case of InfoDisent we apply a few mechanisms to ensure the desired properties. First, we need to be able to disentangle the channel space. To do so we apply the unitary map $U$ in the pixel space. Next, we apply the information bottleneck – for a given channel instead of the average pool where all pixels participate, we use extremely sparse analog where only the value of the highest positive and negative pixels are used. Finally, as is common in interpretable methods we use matrix $A$, but only with nonnegative coefficients to allow only positive reasoning.
The sparse pooling features mechanism
Most interpretable models involve retraining certain parts of the CNN (chen2019looks; rymarczyk2022interpretable), whereas others, like PIP-Net (nauta2023pipnet), retrain the entire CNN. In contrast, our method utilizes a pre-trained CNN or transformer without further modification during training. We first employ a trainable orthogonal transformation $U$ on pixel space to enable the disentanglement of hidden features from feature maps [^1] To parametrize orthogonal maps, we restrict to those with positive determinants and use the formula $U=\exp(W-W^{T})$, where $\exp(\cdot)$ denote the matrix exponential (hall2013lie). We utilize the fact that the space of orthogonal matrices with positive determinants coincide with exponentials of skew-symmetric matrices (shepard2015representation). . To enforce the disentanglement we follow by an introduced sparse analogue of average pooling over channel $K$ given by
$K\to\mathrm{mx\_pool}(K)=\max(\mathrm{ReLU}(K))-\max(\mathrm{ReLU}(-K)).$
Observe that to compute $\mathrm{mx\_pool}(K)$ we need to know only the highest positive and negative pixel values of $K$, contrary to $\mathrm{avg\_pool}$, where all the pixel values are needed. Subsequently, we identify sparse representations (superpixels) within the channels that contribute positively or negatively to predictions. This enables us to generate heatmaps akin to Grad-CAM (selvaraju2017grad) without necessitating a backward model step, as shown in Figure 5(a). Importantly, unlike Grad-CAM, our technique supports the visualization of negative heatmaps, resembling the LRP (bach2015pixel) method that requires a backward pass in a neural network. Our method operates solely during the forward step.
Finally, as is common in XAI models, to allow only positive reasoning we allow the matrix $A$ to have only nonnegative values.
Classification head in InfoDisent
Finally, the classification head in InfoDisent is given by:
  1. 1.
    $I=(I_{rs})_{rs}\to J=(UI_{rs})$, where $U:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is an orthogonal matrix and $I_{rs}$ denotes the pixel value of $I$ with coordinates $r$ and $s$,
  2. 2.
    $J\to v_{J}=\mathrm{mx\_pool\_over\_channels}(J)\in\mathbb{R}^{d}$, where for a given channel $K$ we have
    $\mathrm{mx\_pool}(K)=\max(\mathrm{ReLU}(K))-\max(\mathrm{ReLU}(-K)),$
  3. 3.
    $v_{J}\to w_{J}=Av_{J}$, where $A$ is a matrix with nonnegative coefficients of dimensions $d\times k$
  4. 4.
    $w_{J}\to p_{I}=\mathrm{softmax}(w_{J})$
InfoDisent model
Thus InfoDisent consists of two main components: the frozen CNN or transformer Backbone, and InfoDisent classification head, as illustrated in Figure 2.
The first component, the CNN or transformer Backbone, is a frozen classical pre-trained network up to the final feature maps layer. Importantly, this part of the network and its weights remain unaltered during the training of InfoDisent, ensuring that the learned feature representations are preserved.
The second component involves pooling important features from the final feature maps produced by the Frozen Backbone. This pooling mechanism encourages the network to use information bottlenecks to build maximally informative channels independent of each other. This leads to constructing prototypical channels, which can be easily interpreted. The final element is the fully connected layer, which is distinguished by its positive weight values. This restriction on positivity, except biases handled as in conventional linear layers, ensures that the information about the positive or negative contributions of selected features from the previous part of the model is preserved. This constraint is beneficial for the interpretability of the model’s predictions, as it clarifies the contribution of each feature to the final output.
Observe, that contrary to some Post-hoc methods InfoDisent need training of the model on the whole dataset.
Understanding the classification decisions
Prototypes in InfoDisent
The crucial consequence of InfoDisent that is disentangles the channels making them interpretable. Thus, similarly to PiPNet (nauta2023pipnet), we identify channels as prototypes. To illustrate a given prototype channel, we present five images from the training dataset on which the activation of the channel is the greatest, see Figure 3, where we present consecutive prototypes for the pre-trained ResNet-50 model. This follows from the fact that for better interpretability of model decisions, it is beneficial for humans to be presented from 4 to 9 concepts (rymarczyk2022interpretable). Formally, similarly as in prototypical models, as the prototypical part, we understand the part of the image corresponding to pixels in feature space with maximal activation, marked by the yellow box in  Figure 3. Observe that the presented prototypes seem consistent with each other, and could be well interpreted.
Figure 3: The image shows prototypes from channels 689 to 692 in a trained ResNet-50 on the ImageNet. Each row displays the 5 most significant patches from a single prototypical channel. The prototype’s activations are highlighted by yellow boxes.
Understanding the Decision for a Given Image by Prototypes
[…]
The model’s proposed prototypes are easy to interpret. Moreover, unlike current state-of-the-art prototype methods, our model excels at interpreting images from the ImageNet dataset. More examples from various datasets and models are presented in the Supplementary Materials.
Figure 4: Exemplary explanation (hen) for ResNet-50 backbone provided by InfoDisent in a form of prototypical parts.
(a) The top part shows a heatmap with positive (red) and negative (blue) sparse representations, each channel displaying one positive and one negative superpixel. These sparse representations are then aggregated and rescaled to the image size in the lower part.
(b) Example compares visual explanations from InfoDisent, Grad-CAM, and LRP. The left shows input images, with the following columns displaying explanations from these methods applied to ResNet and DenseNet.
Figure 5: The image demonstrates how to analyze and visualize decisions made by InfoDisent.
Heatmaps
Our approach, which relies on representation channels, enables us to easily generate heatmaps similar to those produced by the Grad-CAM method, see Figure 5(a). To do this we accumulate the activations of all prototypes (both positive and negative ones) over all channels. Since in InfoDisent we use information bottleneck, we obtain more localized results than other standard approaches.
Observe that, unlike Grad-CAM, our heatmaps also illustrate negative activations, akin to the Layer-Wise Relevance Propagation (LRP) method. While LRP can effectively highlight both positive and negative contributions to the model’s decision, it often faces challenges such as complexity in implementation and sensitivity to model architecture variations. Our method addresses these issues by providing clear and interpretable visualizations. Sample results using this explanation method are presented in Figure 24.
Figure 6: Prototypes of two classes (from top to bottom): Half Track and Thatch from the ImageNet dataset. Each row displays the 5 most significant prototypes for a specific channel, with the channel numbers listed on the right. The order of the prototypes within each row reflects their importance for explaining the given class. In the Half Track class, the prototypes from channel 1098 are the most crucial, followed by those from channels 1231 and 1348. These prototypes effectively explain the Half Track class, as they highlight elements of a tank (from channel 1098), a car, and the possible presence of soldiers on board.
Understanding the Decision Behind Class
We examine the model’s decision-making process on a per-class basis by utilizing prototypes. To identify prototypes for a given class, we focus on key channels that are prominently activated across all test set images belonging to that class. These key channels are selected based on their consistent presence and strong activation in images of the same class. Once we have identified these crucial channels, we visualize the prototypes as previously described, providing a clear representation of what the model deems important for that particular class.
Figure 6 illustrates the prototypes for selected classes from the ImageNet dataset. Each prototype captures essential features such as material types, structural elements, or specific textures that are characteristic of the class.
Table 1: Accuracy comparison of interpretability models using standard CNN architectures (utilized in explainable models) trained on cropped bird images of CUB-200-2011, and Stanford Cars (Cars). Our approach demonstrates superior performance across nearly all the datasets and models considered. For each dataset and backbone, we boldface the best result in the class of interpretable models.
Model Dataset
CUB-200-2011 Cars
ResNet-34 ResNet-34 82.4% 92.6%
[Uncaptioned image] InfoDisent (ours) 83.5% 92.8%
[…]
ProtoPShare 74.7% 86.4%
ProtoPool 80.3% 89.3%
ST-ProtoPNet 83.5% 91.4%
TesNet 82.7% 90.9%
ResNet-50 ResNet-50 83.2% 93.1%
[Uncaptioned image] InfoDisent (ours) 83.0% 92.9%
[…]
ProtoTree 86.6%
PIP-Net 82.0% 86.5%
DenseNet-121 DenseNet-121 81.8% 92.1%
[Uncaptioned image] InfoDisent (ours) 82.6% 92.7%
[…]
ProtoPShare 74.7% 84.8%
ProtoPool 73.6% 86.4%
ST-ProtoPNet 85.4% 92.3%
TesNet 84.8% 92.0%
ConvNeXt ConvNeXt-Tiny 83.8% 91.0%
[Uncaptioned image] InfoDisent (ours) 84.1% 90.2%
[…]
Numerical Experiments
This section outlines the experiments conducted to compare our approach with current state-of-the-art methods, highlighting the most significant results. Our experiments utilized a variety of datasets, including well-established benchmarks for evaluating the explainability of artificial models: The Caltech-UCSD Birds-200-2011 (CUB-200-2011) (wah2011caltech), Stanford Cars (krause20133d), and Stanford Dogs (khosla2011novel). Additionally, we incorporated the full ImageNet (russakovsky2015imagenet) dataset for the first time in the class of prototypical networks.
Additional results, including ablation study and analysis of inter-channel correlation, and details concerning those experiments are provided in the Supplementary Materials.
Table 2: Classification accuracy on full CUB-200-2011, and Stanford Dogs datasets by competing approaches using different CNN backbones. For each dataset and backbone, we boldface the best result in the class of interpretable models.
[Uncaptioned image][Uncaptioned image]
Figure 10: Density estimate of the number of significant channels for each class and image in the CUB-200-2011 test set using the ResNet-34 network. InfoDisent uses significantly fewer channels, particularly evident in the middle image, which shows the density estimation of the number of significant channels for all classes.
Figure 11: Density estimate of the number of significant channels for each class and image in the CUB-200-2011 test set using various networks.
Figure 12: Density estimate of the number of significant channels for each class and image in the Stanford Cars test set using various networks.
Appendix D Additional results
In this section, we provide additional results to supplement and expand upon the findings presented in the main part of the paper. We delve deeper into the behavior of our model across various experiments and datasets, offering a more comprehensive analysis.
Ablation study
In this part, we present additional results from a series of analyses investigating the significance of the number of channels on model predictions across various datasets and models. The results of these analyses are illustrated in Understanding the Decision Behind Class, Understanding the Decision Behind Class, Understanding the Decision Behind Class and Understanding the Decision Behind Class. Recall, that InfoDisent model organizes information into channels with sparse representations, which are later utilized in the model’s prediction process. We specifically examine how the number of channels influences the model’s predictions by determining the minimum number of channels required to account for at least 95% of the information used in the model’s predictions.
Formally, if
$\text{logits}=\sum_{i=0}^{N}a_{ki}v_{i}+b_{k},$
where $N$ is the total number of channels, $k$ represents the image class, and $a_{ki},v_{i},b_{k}\in\mathbb{R}$, then for each image from class $k$, we identify the smallest number $n$ of channels such that
$\frac{\sum_{i\in I_{k}}|a_{ki}v_{i}|}{\sum_{i=0}^{N}|a_{ki}v_{i}|}\geq 0.95,$
where $I_{k}$ is the set of indexes of the $n\leq N$ largest values of $|a_{ki}v_{i}|$.
This analysis highlights the most critical channels contributing to the model’s decisions, providing deeper insights into the model’s interpretability and efficiency. Note that the InfoDisent approach consistently utilizes significantly fewer channels, a trend observed across all models and datasets analyzed.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 13: Density estimate of the number of significant channels for each class and image in the ImageNet test set using various networks.
Figure 14: Channel activations before the final linear layer for a randomly selected image processed by different models trained on the CUB-200-2011 dataset are shown. The results are displayed in three groups, each containing two graphs. Model names are listed on the right side of the graphs. Unlike the baseline models, our network activates significantly fewer channels while still maintaining strong performance.
Understanding the Decision Behind Class and Understanding the Decision Behind Class also present the channel values before the final linear layer in our model for randomly selected images from various datasets and models. As evident from the images, our model utilizes a significantly smaller number of channels in its predictions compared to the baseline models.
Figure 15: Channel activations before the final linear layer for a random image processed through various ConvNeXt, Vision Transformer, and Swin Transformer models trained on ImageNet. The results are displayed in three groups, each containing two graphs. Model names are listed on the right side of the graphs. InfoDisent activates significantly fewer channels compared to the baseline models while maintaining high effectiveness.
Explaining Classification Decision for a Given Image by Prototypes
InfoDisent employs prototypes, similar to the approach used in PiPNet [nauta2023pipnet], to explain individual decisions made on an image. Below, we describe the process for identifying prototypes for a given input image. An average pooling operation makes the aggregation of information from individual channels. The outcome of this operation is a scalar for each channel $K$, computed as
$\mathrm{mx\_pool}(K)=\max(\mathrm{ReLU}(K))-\max(\mathrm{ReLU}(-K))$
(as explained in the main paper). This dense representation, formed by aggregating all the channels, is then processed by a linear layer that outputs logits, see Figure 2. The logits can be represented in a format similar to the output of a convolutional layer, as illustrated by $V_{+}$ and $V_{-}$ in Figure 2. This approach maintains the pictorial structure of the logits, allowing us to extract individual channels.
Note that each channel in InfoDisent model can contain only two possible values (refer to Figure 5a in the main paper, where these values are depicted as red or blue areas within the channels). To identify the prototype, i.e., the relevant channel, we focus solely on the positive values within the channels (represented by red areas in Figure 5a). These positive values indicate the significant part of the channel/prototype (marked by the yellow frame on the prototype) and define the channel’s importance for the model prediction (since the linear layer uses only nonnegative coefficients). As demonstrated in Understanding the Decision Behind Class and Understanding the Decision Behind Class, the number of such channels is limited (alternatively, we could also focus on channels with the strongest values). Once we have identified the important channels (by knowing their indices), we represent each channel using prototypes. Prototypes are images from the training set that exhibit the five strongest positive values for a given channel. Example results illustrating how model predictions are explained using prototypes are shown in Figures 16, 17, 18, 19 and 20 (which are after the references).
Figures 16, 17 and 18 showcase the performance of our prototype models on standard benchmarks: CUB-200-2011, Stanford Cars, and Stanford Dogs. Figure 16 demonstrates the model’s ability to focus on distinctive features like the Scissor-tailed Flycatcher’s elongated tail feathers, underwing yellow coloration, or the Red-legged Kittiwake’s red feet.
Similarly, Figure 17 highlights the model’s capacity to identify key vehicle components. For instance, it accurately pinpoints the fender, bumper, and body of a Jeep Wrangler SUV 2012, and the characteristic stripes on a Ford GT Coupe 2006.
Complementing our experiments with widely used CNN models, we investigated the performance of transformer architectures, specifically VisionTransformer (ViT-B/16) [dosovitskiy2020image], SwinTransformer (Swin-S) [liu2022swin]. Their results are visualized in Figures 19 and 20. Our analysis reveals that transformer models concentrate on smaller image regions than CNNs. Nevertheless, both model types generate interpretable prototypes that offer insights into the input image content.
Decision Behind Class
To delve deeper into the model’s decision-making process, we employ a prototype-based analysis at the class level. For each class, we identify key channels that consistently exhibit strong activation across the entire test set. These channels, indicative of the model’s focus on specific visual features, are selected based on their prominence and reliability in representing the class. By visualizing these key channels as prototypes, as demonstrated in our previous image analysis, we obtain a clear representation of the model’s class-specific decision criteria.
Figures 21, 22 and 23 present the results of our prototype analysis for selected classes in the CUB-200-2011 and Stanford Cars datasets, using both InfoDisent(ResNet-50) and transformer models. This visualization allows for a granular understanding of how these models differentiate between various classes, highlighting the underlying patterns recognized by the network.
Figure 16: Example prototypes generated by InfoDisent(ResNet-50) models for random input images from the test set of the CUB-200-2011 dataset. The display includes 8 input images, each with a corresponding column where yellow boxes highlight specific regions, followed by the prototypes (images on the right). Each row represents prototypes from a different channel, with the channel index on the right. Observe that the prototypes identified by the model effectively capture distinct parts of the body in the images.
Figure 17: To demonstrate InfoDisent(ResNet-50) model’s capability, we visualize prototypes generated for arbitrary test images from the Stanford Cars dataset. Each image is paired with highlighted areas and its associated prototypes, grouped by channel. The results indicate that the model has learned to extract meaningful features representing different vehicle parts.
Figure 18: Visualized prototypes from the Stanford Dogs dataset. Each row highlights prototypes from a specific channel, focusing on different dog features such as ears, nose, and fur.
Figure 19: This figure showcases the prototypical explanations provided by InfoDisent(ViT-B/16) model. We visualize prototypes for arbitrary ImageNet images, highlighting relevant regions and their corresponding channel-specific prototypes.
Figure 20: The figure demonstrates the prototypical explainability of InfoDisent(Swin-S) model by visualizing prototypes generated for arbitrary ImageNet images, highlighting relevant image regions and their corresponding channel-specific prototypes.
Figure 21: Prototypes of sample classes from InfoDisent(ResNet-50) model trained on the cropped Stanford Cars dataset (left) and the CUB-200-2011 dataset (right). Yellow frames highlight the class prototypes.
Figure 22: Visualizations of InfoDisent(ViT-B/16) model-generated prototypes for selected ImageNet categories.
Figure 23: Representative ImageNet class prototypes produced by the InfoDisent(Swin-S) model.
Heatmaps
Figures 24 and 25 present example heatmaps generated by our model, resembling those produced by Grad-CAM [selvaraju2017grad]. Our method, rooted in representational channels, simplifies heatmap generation by accumulating activations from all prototypes (positive and negative) across all channels. Leveraging the information bottleneck principle, our approach yields more focused heatmaps compared to traditional methods.
More details on user study
Each worker was paid €2.00 for completing a short 20-question survey. The survey questions were randomly composed, so the specific questions differed between participants. The participants were gender-balanced and ranged in age from 18 to 60. They were given 30 minutes to complete the survey.
To ensure data quality, we excluded responses where users selected the same answer for all questions. Surveys were repeated until we obtained 60 valid responses. Figure 26 and Figure 27 illustrate example questions used in both user studies.
Before starting the survey, participants were provided with an example and detailed instructions to familiarize them with the study setup, including the explanation composition and visualization. The distribution of answers is summarized in Tables 6 and Table 7.
Table 6: Distribution of answers in user study on user confidence in model’s prediction based on explanation.
Dataset Fairly … correct Somewhat … correct Somewhat … incorrect Fairly … incorrect
CUB 585 316 208 91
ImageNet 449 248 189 314
Table 7: Distribution of answers in user study on prototypical part disambiguity.
Dataset Correct Incorrect
CUB 776 424
ImageNet 712 488
Is InfoDisent reducing channel correlation?
To assess whether our model reduces channel correlation, we computed the RV coefficient, a standard measure of linear correlation between vectors. In 14 experiments (across 3 datasets and 7 backbones), our model outperformed the baseline in 9 cases (smaller RV coefficients), performed worse in 2 cases, and showed equivalent results in 3 cases as presented in Table 8. Despite not having a loss term explicitly aimed at reducing the RV coefficient, our model generally demonstrates either smaller or equal RV coefficients compared to the baseline.
Table 8: RV correlation between channels for InfoDisent and baseline models.
Dataset Model InfoDisent Baseline
Cars ConvNeXt Tiny 6.7 6.9
DenseNet121 6.2 6.9
ResNet34 6.0 7.1
ResNet50 6.2 7.6
CUB-200-2011 ConvNeXt Tiny 5.5 5.5
DenseNet121 7.2 7.9
ResNet34 6.5 7.2
ResNet50 4.9 7.8
ImageNet ConvNeXt Large 8.7 3.0
DenseNet121 8.4 4.5
ResNet34 5.2 6.0
ResNet50 3.0 6.5
SwinV2-S 2.5 2.5
ViT-B-16 9.9 9.9
Figure 24: Comparison of example heat maps generated by InfoDisent proposed model and competing approaches. Our technique produces more focused heat maps by leveraging representational channels and the information bottleneck principle, outperforming traditional methods like Grad-CAM.
Figure 25: Example heatmaps generated by InfoDisent, our proposed method, demonstrating activation regions on sample photos from the ImageNet test set.
Figure 26: An exemplary question from the user study on user confidence.
Figure 27: An exemplary question from the user study on disambiguity of prototypical parts.